The circle x^2 + y^2 - 2x - 3 = 0 is stretched horizontally by a factor of 2 about x = 0 to obtain an ellipse. what is the equation of this ellipse in general form
Answer by Edwin McCravy(20059)



To stretch horizontally by a factor of k, the rule is:

Replace x by x/k

So we replace x by x/2 in

        xイ + yイ - 2x - 3 = 0

(x/2)イ + yイ - 2(x/2) - 3 = 0

       xイ/4 + yイ - x - 3 = 0

Cl;earing of fractions:

      xイ + 4yイ - 4x - 12 = 0

That's all you were supposed to do.  However if you
place the circle and the ellipse in standard form
you get:

(x - 1)イ + (y - 0)イ = 2イ

which is this graph:



And the ellipse has the standard equation

(x - 2)イ   (y - 0)イ
覧覧覧覧 + 覧覧覧覧 = 1
   4イ         2イ

and its graph is:



Edwin